OpenStudy (anonymous):

find the function values. g(-1)=^4 square root of t-8. I got 3. g(6569)= same problem. I got 81. or should it be 3? g(9)= same problem. I got 1. I just want to make sure I am on the right track.

OpenStudy (dumbcow):

\[\sqrt[4]{t-8} = 3\] when t = -1 is that what you are saying ?

OpenStudy (caozeyuan):

well if you can write you function more clearly, I may be able to help you. I recommend you to use the equation button below

OpenStudy (caozeyuan):

but when t=-1 the value to that function above should be an imaginary number

OpenStudy (anonymous):

\[g(-1)=\sqrt[4]{t-8}\]

OpenStudy (caozeyuan):

so what are you trying to find?

OpenStudy (caozeyuan):

or you mean g(x)\[=\sqrt[4]{t-8}\]

OpenStudy (caozeyuan):

and set t= -1?

OpenStudy (anonymous):

the g(t) and then i am given values for t

OpenStudy (caozeyuan):

ok so just plug in your number into that equation

OpenStudy (anonymous):

for g(-1) i got an answer of 3. for g(6569) i got an answer of 81. for g(9) i got an answer of 1.

OpenStudy (anonymous):

but what do i do with the 4 at the beginning?

OpenStudy (caozeyuan):

well if you plug in -1 into the equation above, thats the 4th root of (-1-8) so the 4th root of -9 is the sqrt of 3i, so\[\sqrt{3i}\] is your answer